# The contact process on dynamical random trees with degree dependence

### v1

@@ 7.4 Transmitting the infection from one star to another
In addition, by definition of $r_{N}$, we have $cN\mu^{r-1}_{L}\mathbb{P}(\zeta=N)\geq 1$, which implies $H(x)\geq x$ in a neighborhood of 0. Thus appealing to Lemma 3.4 in [18], we conclude that
$\liminf_{t\to\infty}g(t)>0,$
and thus
$\liminf_{t\to\infty}\mathbb{P}^{\{\rho\}}\big{(}\rho\in\mathbf{C}_{t}\big{)}\geq\mathbb{P}(\zeta=N)\liminf_{t\to\infty}g(t)>0$
which in turn implies that $\lambda_{2}=0$.
For the second part, we fix $\eta>0$ and $\alpha<1$ and assume that Condition (3.8) holds. Now, for $N\in\mathbb{N}$, we choose $\lambda_{N}=(3T_{N})^{-1}$ and $r=r_{N}$ as in (7.30). By Lemma 7.13, we know that for some large $N$ the term $r_{N}$ satisfies Condition ($\star$7.11). Therefore, using the same arguments as before we can conclude that for this choice of the infection rate,
$\liminf_{t\to\infty}\mathbb{P}^{\{\rho\}}\big{(}\rho\in\mathbf{C}_{t}\big{)}>0,$
which allows us to conclude that $\lambda_{2}\leq\lambda_{N}<\infty$. ∎
Acknowledgement. MS was supported by the LOEWE programme of the state of Hessen (CMMS) in the course of this project.

### v2

@@ 7.4 Transmitting the infection from one star to another
In addition, by definition of $r_{N}$, we have $cN\mu^{r-1}_{L}\mathbb{P}(\zeta=N)\geq 1$, which implies $H(x)\geq x$ in a neighborhood of 0. Thus appealing to Lemma 2.4 in [18], we conclude that $\liminf_{t\to\infty}g(t)>0,$ and thus
$\liminf_{t\to\infty}\mathbb{P}^{\{\rho\}}\big(\rho\in\mathbf{C}_{t}\big)\geq\mathbb{P}(\zeta=N)\liminf_{t\to\infty}g(t)>0.$
It follows from the reverse Fatou lemma (which applies as indicator functions are bounded by $1$) that
$\mathbb{P}^{\{\rho\}}\big(\forall s\geq 0\,\exists t\geq s\,:\,\rho\in\mathbf{C}_{t}\big)=\mathbb{E}\Big[\limsup_{t\rightarrow\infty}\mathds{1}_{\{\rho\in\mathbf{C}_{t}\}}\Big]\geq\limsup_{t\to\infty}\mathbb{P}^{\{\rho\}}\big(\rho\in\mathbf{C}_{t}\big)>0.$
Thus, as $\lambda>0$ was arbitrary, it follows from Lemma 5.2 that $\lambda_{2}=0$.
For the second part, we fix $\eta>0$ and $\alpha<1$ and assume that Condition (3.8) holds. Now, for $N\in\mathbb{N}$, we choose $\lambda_{N}=(3T_{N})^{-1}$ and $r=r_{N}$ as in (7.29). By Lemma 7.14, we know that for some large $N$ the term $r_{N}$ satisfies Condition ($\star$7.12). Therefore, using the same arguments as before we can conclude that for this choice of the infection rate, $\liminf_{t\to\infty}\mathbb{P}^{\{\rho\}}\big(\rho\in\mathbf{C}_{t}\big)>0$, which allows us to conclude in the same way as above that $\lambda_{2}\leq\lambda_{N}<\infty$. ∎
Acknowledgement. The authors would like to thank the anonymous referee for their detailed comments and suggestions which helped us to improve the article. MS was supported by the LOEWE programme of the state of Hessen (CMMS) in the course of this project.